Suppose you double the lengths of the sides of a rectangle.
What is the relationship between the perimeter of the new rectangle and the perimeter of the original rectangle?

Answer by checkley77(12844)

Let L & W be the original sides.
2L+2W=perimeter.
Let 2L & 2W be the expanded sides.
2*2L+2*2W=perimeter
4L+4W=perimeter for the expanded redtangle.
The perimeter is doubled.